# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 9.3.8 A case study of larch moth interactions
Plant’s quality plays an important role in the population dynamics of consumers [127, 128]. Din and Khan explored the dynamics between budmoths and the quality of larch trees located in the Swiss Alps [80]. They proposed an improved discrete-time consumer-resource model to capture the interactions between the plant quality index $x$ and the moth population density $y$ defined as
@@ 11 Conclusion
In this article, we presented a comprehensive and systematic review of the literature concerning the stability of ecological systems. Notably, we inclusively surveyed various stability notions that arise in ecological systems, encompassing linear stability, sign stability, diagonal stability, D-stability, total stability, structural stability, and higher-order stability.
As mentioned earlier, empirically applying this theory to real ecological systems poses a significant challenge. First, most of the stability criteria are derived from the random matrix theory, which cannot accurately capture real ecological systems. For instance, the food web topology studies have demonstrated the non-randomness of ecological systems, rendering to incorporate more realistic network structures [150]. Second, the dynamics governing real ecological systems could not adhere to the GLV model, or even unknown in many cases. Thus, the stability analysis of real ecological systems may not align with the predictions of the theory.

### 07 Feb 2024 v2

@@ 10.3.8 A case study of larch moth interactions
Plant’s quality plays an important role in the population dynamics of consumers jang2009dynamics ; jang2009models . Din and Khan explored the dynamics between budmoths and the quality of larch trees located in the Swiss Alps din2021discrete . They proposed an improved discrete-time consumer-resource model to capture the interactions between the plant quality index $x$ and the moth population density $y$ defined as
@@ 11 Discussion
Sector stability analysis is particularly useful when dealing with semi-feasible equilibrium points, which are prevalent in modeling ecological systems. Local sector stability can be readily assessed, but a definitive algorithmic approach for verifying global sector stability remains elusive. For example, how can we efficiently compute the crucial constant matrix G defined in Section 8? Moreover, it will be worthwhile to explore simplified conditions on the sector stability of other nonlinear ecological models such as the GLV model with higher-order interactions and MacArthur’s consumer-resource model.
@@ 12 Conclusion
In this article, we presented a comprehensive and systematic review of the literature concerning the stability of ecological systems. Notably, we inclusively surveyed various stability notions that arise in ecological systems, encompassing linear stability, sign stability, diagonal stability, D-stability, total stability, sector stability, structural stability, and higher-order stability.
As mentioned earlier, empirically applying this theory to real ecological systems poses a significant challenge. First, most of the stability criteria are derived from the random matrix theory, which cannot accurately capture real ecological systems. For instance, the food web topology studies have demonstrated the non-randomness of ecological systems, rendering to incorporate more realistic network structures landi2018complexity . Second, the dynamics governing real ecological systems could not adhere to the GLV model, or even unknown in many cases. Thus, the stability analysis of real ecological systems may not align with the predictions of the theory.
@@
Acknowledgements
This work was supported by the National Institutes of Health R01AI141529.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT in order to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.